# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v2

@@ VII Sim-to-Real Gap Reduction Techniques
sim-to-real is notoriously difficult. This work handles two kinds of sim-to-real gaps: the physical dynamics and the visual gaps. For brevity, we discuss a subset of the techniques which are crucial in addressing sim-to-real and for successful real-world deployment. More techniques can be found in Appendix D.
VII-A Dynamics Gap Reduction
The dynamics gap is essentially caused by misaligned transition functions $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ between the simulated and real worlds. This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance. In fact, even minor covariate shifts can result in task failure if the task requires fine motor control, e.g., grasping a small object. Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real. In particular, we use the following key techniques for addressing the issue: 1) arm joint tracking error minimization, 2) domain randomization, and 3) object perturbations.
Minimizing Arm Tracking Error via PID Control.
We used PID controller for the robot arm. As opposed to PD controllers which will always have gravity-induced tracking errors, the integral term of PID controllers allow for more robust pose-agnostic tracking given that the joint delta targets are sufficiently small for particular operating frequency. This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification. More details can be found in Appendix D-A.
Arm Control Perturbation.
Random structured noise is added to the arm control signal to make the learned policy to be robust w.r.t. different sources of control noises arising when transferring to real.
Arm Mount Perturbation.
We randomly perturb the arm mount position and yaw to make the learned policy robust to the sim-to-realgaps that could arise due to the actual mounting accuracy, torso height variations etc..
Object Perturbations.
The final strategy we employ are object perturbations. Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies. For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base. This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off. Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution. Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robust ness of the policy (as the state coverage is enlarged).
Stationary Manipulations.
We make the quadruped stand as still as possible when manipulating objects, i.e. during grasp or drop into. This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed. To achieve this, when training low level policies to follow locomotion speed commands, we also make the quadruped enter a stationary mode with a balanced stance when the speed command is close to zero. Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers and arm motors.
VII-B Visual Gap Reduction
@@
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by an RGB distribution mismatch between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
Image Domain Randomization.
Random Textures.
We spawn the robot on an empty ground and randomize the ground texture. The texture can be arbitrary and generally have no specific semantics, as long as they provide enough visual distraction to the student. This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors. When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
Random Spatial Augmentation.
Random Background Objects.
To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position. There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere), and 2) Google scanned objects [10]. We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects. It could collide with an background object during training with some penalty reward. When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
Refer to caption Refer to caption
Figure 6: An example scene populated with random background objects. Left: close-up view; Right: birds-eye view.
@@ VIII Physical System, Experiments and Results
In this section, we briefly go over the robot system and the task used to evaluate SLIM.
@@ VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig. 7, serving as the robot’s sole RGB vision feed. The depth stream is not used in this work. Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD. With the non-edu version, total cost is even lower ($\sim$6K USD) and reflects a bit more closely the cost of mass produced parts..
@@
VIII-B Baselines
We compare SLIM with a number of baselines below:
  • No Arm Retract: removing the arm retract (raise) reward in RL training, which is used to encourage the robot’s arm to raise up after grasping the object (introduced in Section VI);
  • No Perturb: removing arm control randomization, arm mount perturbation and object perturbation (introduced in Section VII-A)
  • No Visual Aug: no background objects and no random spatial augmentation based visual augmentation (introduced in Section VII-B)
  • Distillation Only: agent training without using the RL loss (introduced in Section V), keeping only the distillation and representation loss.
  • Human Teleop: an expert human trained for half an hour in teleoperating the robot with locomotion and arm end-effector commands observing only the view of the wrist camera. While not a baseline per se, human teleoperation provides a valuable reference point. See Appendix B-D for more details.
We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e. RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. It is excluded from the tables.
VIII-C Evaluation Protocol
We use the Standard object spatial layout as shown in Figure 8 in a real-world scene Lobby. By doing so, it is easier to maintain repeatability across all methods and seeds. Besides the Standard in lobby, we also evaluate the robot in several other scenes with other layout. For task objects, we use cubes and baskets of different colors. More details on these physical objects are detailed in Appendix F-A.
For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene setup (i.e. different positioning and colors of objects) according to the evaluation protocol. For the detailed protocol, please refer to Appendix F-B.
TABLE I: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
Method Fully Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Autonomous Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
@@ VIII-D Results and Analysis
Figure 8: Scene Spatial Layouts. Standard: we put the robot in the center of a square space and objects including objects to be grasped and containers to drop the graspable objects into on the four corners. The instructions and object placements vary across episodes. Cluttered: objects are placed close to each other in front and behind the robot, leading to OOD spatial layouts. The instructions and object placements vary across episodes.
We conduct experiments on the baselines and the proposed method, and measure their success rates in the real world. The results are summarized in Table I.
We run the whole training pipeline with three different random seeds to train low-level, the teacher and the student, and measure the performance of the final student policy in real, in terms of accuracy and efficiency:
  1. 1.
    Cumulative Subtask Success Rate: the success rate up to the specified subtask from the beginning of the episode.
  2. 2.
    Episode Time: the time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$s. For failed episodes, we use $t_{\rm max}$ as episode time.
Finally, we report the mean and standard deviation across the seeds as the final aggregated metric value in Table I.
Table I shows how important each of the interventions used contribute to the final successes. No Arm Retract baseline performs well for the initial two subtasks of Search and MoveTo, but not for the later subtasks. Even though half of the times the robot can move to the basket carrying the cube (a cumulative success rate of $28.3\%$), it only successfully drops the object into the basket less than one fifth of the cases (a cumulative success rate of $5.0\%$). The main reason is that without arm retract reward, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm. This leads to more arm shake, cube slipping out of the fingers, motor failures, and makes it harder for the arm to raise up to search for the basket and to hover above the basket to complete DropInto. No Perturb baseline has the lowest grasping success rate. Without perturbations (object, arm mount and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot. This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen. No Visual Aug’s performance on the initial Search subtask is lower than the that of the above baselines. The main reason is that without visual augmentations, it can be distracted by background objects irrelevant to the task, i.e. false positive detection of the target. This shows that visual augmentations are important for improving the robustness of the robot’s vision module. Distillation Only baseline simply imitates from the highly successful teacher policy, and can indeed achieve a relatively high success rate. However, its performance is not consistent across random seeds. Our full method (SLIM) uses distillation-guided RL, and results in consistently high success rates across seeds. It shows that distillation loss alone is not enough for robust policy learning. Additionally, task completion is also slower than the full method which additionally uses RL loss to optimize the policy.
Our full method SLIM achieves the highest subtask and full task success rate among all the methods, and is also the most efficient, taking only 43.8 seconds to complete the full task on average.
@@
Figure 9: Real-world Performance of SLIM across Scene Variations. Left: images showing the differences for evaluation. Right: the full task success rate across difference scenes.
@@ VIII-E Generalization to Different Real-world Scenes
In addition to the standard scene used in Table I, we further run real-world evaluations of SLIM under more scene variations (c.f. Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby. For the first three scenes, we use the standard spatial layout (c.f. Figure 8). For Lobby, we use the Cluttered spatial layout, which is different from the training (Standard) layout. Results as summarized in Figure 9 are close to the Standard layout or a bit lower due to being out of distribution. This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, background, and floor materials and textures.
@@
    task objects are randomly scattered on the ground (e.g., last two rows in Figure 1),
@@
    distractor objects are present (e.g., last row in Figure 1),
@@
    human interrupts the task progress, or
@@
    objects to grasp have slightly different shapes.
We refer the reader to our demo videos for more emergent generalization behaviors. In general, the robot is quite robust to the real-world scene complexity.
VIII-F Comparison with Simulation Performance
Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variability in the RL training of the teacher policies. Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well.
Figure 10 shows simulation performance with the real performance for our methods, so we know more clearly how each method helps the training process.
The Sim success rate for each method denotes the average full task success rate over 200 episodes run in simulation, averaged across the three random seeds. The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy. The Real success rate denotes the full task success rate of the policy in the real-world, following the Standard spatial layout as shown in Figure 8.
Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasp, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up, creating an exploration bottleneck for RL training. Real world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: The tip of the arm tends to shake more due to motor backlash, hard to model in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overload. For No Perturb, while one might expect perturbations to complicate RL learning, their inclusion results in a much higher performance even in simulation. Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups. Without visual augmentations (No Visual Aug), simulation task is easier to solve, thus simulation success is artificially high, but real world performance suffers when compared to the full method. RL training loss in the full SLIM method increases success in simulation (over the imitation only Distillation Only) and achieves the highest success rate in real. It’s interesting to note that the RL loss seems to help the No Visual Aug baseline to achieve slightly better real world success than Distillation Only.
@@ VIII-F Comparison with Simulation Performance
Figure 10: Full Task Success Rates of Different Methods (averaged across three seeds) in Simulation and Real.
[…]
Figure 11: Policy Trajectory Visualization. For visualization purpose, we removed background objects. We also visualize the ee trajectories. Trajectories corresponding to different subtasks are rendered with different colors. The ego-centric camera view is shown on the top-right of each image. The red colored arrows denote the direction of movement.
VIII-G Policy Behavior Visualization
The behavior of the policy is shown in Figure 11. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Figure 11-Search). Once the target object is in view, the robot will move towards it (Figure 11-MoveTo). Once the robot moves close to the target object, it will start to grasp (Figure 11-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE trajectory in Figure 11-SearchWithObj. Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
VIII-H Failure Modes
We also recorded common failure cases for SLIM. In one case, when grasping the quadruped stopped too late and kicked the cube far away. In the other case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. DropInto sometimes fails because of either early dropping right outside of the basket, or the quadruped kicking the basket away right before dropping when backing towards the basket. In rare cases where the gripper holds only the upper half of a cube, the robot tends to get stuck at hovering over the basket while holding the cube, because the cube is completely blocked by the gripper fingers in the wrist camera view. See our failure demo videos for more details.

### v3

@@ VII Sim-to-Real Gap Reduction Techniques
In this section, we discuss crucial techniques for addressing sim-to-real gaps and achieving successful real-world deployment.
The dynamics sim-to-real gap is the mismatch between the transition function of the simulation versus that of the real world. It can be due to the mismatch in the physical properties of objects and motors, or by things that are not properly simulated, such as friction and backlash. This difference in dynamics can cause policies trained purely in simulation to fail when deployed in the real world, especially if the task requires accurate motor control, such as grasping a small object.
@@
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by an RGB distribution mismatch between simulated and real-world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios.
For brevity, we list techniques we deemed most relevant in Table I. Comprehensive details can be found in Appendix E.
@@ VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX-250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount. The RGB stream of D435 is used as the robot’s sole vision feed (Fig. 6). Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Finally, the base of the WidowX is raised a few inches to allow for easy access to the onboard Raspberry PI’s ports. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD. With the non-edu version, total cost is even lower ($\sim$6K USD) and reflects a bit more closely the cost of mass produced parts..
@@
TABLE II: Dimension values for the observations and actions of our robot. For the meanings of these notations, we refer the reader to Section III.
$L$ $H$ $W$ $N$ $D$ $K$ $Q_{m}$ $M$ $z$ $I$ $J$
$100$ $90$ $160$ $5$ $66$ $7$ $19$ $6$ $0.05$ $6$ $12$
@@
Given our hardware configuration, the dimension values from Section III are summarized in Table II.
VIII-B Ablations and Baselines
We compare SLIM against the following baselines, where the first four are ablations and the last is a human teleoperator as a reference:
  • No Arm Retract: Remove the arm retract (raise) reward in RL training (Section VI).
  • No Perturbation: Remove arm control randomization, arm mount perturbation, and object perturbation (Section E-A).
  • No Visual Aug: Remove random background objects and spatial visual augmentations (Section E-B).
  • Distillation Only: Train student without the RL loss, keeping only the distillation and representation loss (Section V).
  • Human Teleop: An expert human teleoperator that provides locomotion commands and delta EE pose to the robot via a joystick. Delta EE is handled via IK. The teleoperator shares the same observation space as the policy, i.e., observing solely through the wrist camera stream, and is allowed to practice on the system for an hour. While not a baseline intrinsically, human teleoperation provides a valuable reference point. See Appendix I for more details.
We also tried another baseline No Distillation, which trains the student policy directly via RL without distillation from the teacher policy (i.e., RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. As such, it is excluded from the results.
VIII-C Metrics and Evaluation Protocol
To accurately evaluate policy performance, we run the entire training pipeline across three different random seeds for all methods, training from scratch the low-level, teacher, and student policies. Note that each method for a particular seed shares the same low-level policy. We report the following two key metrics when deploying the final student policy:
  1. 1.
    Cumulative Subtask Success Rate: The success rate up to each subtask from the beginning of the episode.
  2. 2.
    Episode Time: The time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$ s. For failed episodes, we use $t_{\rm max}$ as the episode time.
For the deployment environment, we use the Standard object spatial layout in an indoor Lobby scene, as shown in Fig. 7, to maintain repeatability across all methods and seeds. For task objects, we use cubes and baskets of different colors. For each method of each random seed, we roll out the robot’s policy for 20 episodes with varying object types and positions according to a consistent evaluation protocol, with 360 real-world episodes of experiments. More details on the physical objects and protocol can be found in Appendices G-A and G-B, respectively. In addition to the Standard spatial setup in Lobby, we also evaluate the robot in several other scene and layout combinations in subsequent experiments, with another 40 real-world episodes, leading to a total of 400 real-world episodes of experiments.
TABLE III: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
Method Autonomous Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
@@ VIII-D Results and Analysis
Figure 7: Scene Spatial Layouts in Lobby Environment. Standard: we put the robot in the center of a square space and objects including objects to be grasped and containers to drop the graspable objects into on the four corners. The instructions and object placements vary across episodes. Cluttered: objects are placed close to each other in front and behind the robot, leading to an OOD spatial layouts. The instructions and object placements vary across episodes.
All success rate and task completion results are summarized in Table III. As shown, we can see the negative impact of ablating away each technique on the final task success.
In particular, No Arm Retract performs fairly well up until Grasp, but then soon its performance drops off substantially. In fact, No Arm Retract possesses the lowest cumulative success rates past grasping across the board, with a sharp dropoff of approximately 30% when attempting to search the container. There are several reasons for this dropoff. For one, without the arm retract reward acting as a soft constraint on the arm’s position, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm. This leads to more arm shake, which in turn can result in the grasped cube slipping out of the fingers or even motor failures. Another key reason for search-based failure arises from possible low vantage search strategies, e.g., the camera is pointed toward the ground.
Intuitively, No Perturbation has the lowest grasping success rate. This makes sense as without perturbations of any kind (object, arm mount, and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot. This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen.
Performance of No Visual Aug on each of the Search subtasks suffer as the model sometimes gets distracted, having false positive detections of the target in the background. There’s a particularly large drop of roughly 25% when searching for the basket. This indicates that visual augmentations are important for improving the robustness of the robot’s vision module.
Finally, Distillation Only simply distills the highly successful teacher policy, and can indeed achieve a relatively high success rate. However, its performance is not consistent across random seeds. In comparison, SLIM uses distillation-guided RL, resulting in consistently high success rates across all seeds indicated by the low standard deviations. It shows that distillation loss alone is not enough for robust policy learning. Additionally, task completion is also slower as the RL loss allows the policy to further optimize. Overall, SLIM achieves the highest subtask and full task success rates among all the methods including Human Teleop, and is also the most efficient, taking on average 43.8 seconds to complete the full task, which is about 1.5$\times$ the speed of a teleoperation expert.
We provide some visualization of the policy behaviors and prediction outputs in Appendix H.
@@
Figure 8: The same SLIM Policy across Scene Variations. Left: images showing the differences for evaluation. Right: the average full task success rate across 10 real-world deployment episodes for each of the scenes shown on the left.
@@ VIII-E Generalization to Different Real-world Scenes
In addition to the standard scene used in Table III, we further run real-world evaluations of SLIM under more scene variations (left side of Fig. 8): Outdoor, Carpet Room, Kitchen, and Lobby-Cluttered. For the first three scenes, we use the same Standard spatial layout as before. For the Lobby-Cluttered scene, we use the Cluttered spatial layout (right side of Fig. 7), which is different from the Standard layout used during training. We conduct ten trials for each environment using the same SLIM policy.
All results for these four additional scenes are summarized in the right side fo Fig. 8. For the three novel scenes, we observe that success rates are all fairly close to the results from Table III ($\sim$78%). For the Lobby-Cluttered scene, we can see that the success rate is quite close to the Standard layout and is just a bit lower due to being out of distribution. This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, backgrounds, and floor materials and textures.
@@
    task objects are randomly scattered on the ground (last two rows in Fig. 1),
@@
    distractor objects are present (last row in Fig. 1),
@@
    a human interrupts the task progress (in demo videos), or
@@
    objects to grasp have slightly different shapes (in demo videos).
We refer the reader to our demo videos in the supplementary file for more emergent generalization behaviors.
VIII-F Sim vs. Real Performance
Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variance in the RL training of the teacher policies. Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well. To better understand how each technique impacts sim-to-real gap reduction, we showcase the simulation and real performance difference for all methods in Fig. 9. The sim success rate for each method denotes the average full task success rate over 100 episodes run in simulation, averaged across the three random seeds. The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy. The real success rate denotes the full task success rates from Table III.
Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasping, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up (without explicit rewards encouraging it), creating an exploration bottleneck for RL. Real-world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: the tip of the arm tends to shake more due to motor backlash, which are hard to be modeled precisely in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overloading. For No Perturbation, while one might expect perturbations to complicate RL learning, their inclusion results in much higher performance even in simulation. Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups. Without visual augmentations (No Visual Aug), the simulation task is easier to be solved, and thus simulation success is artificially high, but real-world performance suffers when compared to the full method. RL training loss in the full SLIM method increases success in simulation (over Distillation Only) and achieves the highest success rate in real.
@@
Figure 9: Full Task Sim and Real Success Rate Comparison. All success rates are the averaged success rate across three random seeds.
VIII-G Failure Modes
We also recorded failures for SLIM and report the most common cases here. In one case, the quadruped stopped a bit late during grasping and kicked the cube away. In another case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. DropInto sometimes failed because of either early dropping right outside of the basket, or the quadruped kicking the basket away, right before dropping, when backing towards the basket. In rare cases where the gripper held only the top part of the cube, the gripper fingers completely blocked the cube from the camera view, and the robot got stuck hovering over the basket while holding the cube. See our failure demo videos for more details.